Prove that for every subgroup H of Sn, either all of the permutations in H are even, or exactly half are even.



1
Expert's answer
2021-04-14T00:48:19-0400

Let "a" be the number of odd permutations of H and "b" denotes the number of even permutations. Then let "\{g_{1},g_{2},\cdots, g_{a}\}" denotes the odd permutations. Then we consider ".S=\{g_{1}g_{1},g_{2}g_{1},g_{3}g_{1}\cdots g_{1}g_{a}\}" . By closure property "S\subseteq H." Also "g_{1}g_{i}=g_{1}g_{j}\Rightarrow g_{i}=g_{j}." Hence, "|S|=a." "x,y" odd permutations implies "xy" even. Hence "S" consists of even permutations. Hence "a\leq b [\because S\subseteq H]" . Now even permutations of a subgroup always forms a group since, inverse of a even permutation is even and "x,y \ even \Rightarrow xy \ even." Also "identity" is even permutation. Hence the even permutations are a subgroup of H. Hence by Lagrange, "b|a+b" . This shows, "b|a." So either "a=0," or "a=b [\because b\geq a]." Hence "b=|H| \ [case: a=0]\ or \ b=\frac{|H|}{2} \ [a=b]."